### v1

Abstract
To tackle the crucial problem of crime, evidence at indoor crime scenes must be analyzed before it becomes contaminated or degraded. Here, as an application of artificial intelligence (AI), computer vision, and robotics, we explore how a blimp could be designed as a kind of "floating camera" to drift over and record evidence with minimal disturbance. In particular, rapid prototyping is used to develop a proof-of-concept to gain insight into what such blimps could do, manually piloted or semi-autonomously. As a result, we show the feasibility of attaching various components to an indoor blimp, and confirm our basic premise, that blimps can sense evidence without producing much wind. Some additional suggestions–regarding mapping, sensing, and path-finding–aim to stimulate the flow of ideas for further exploration.
keywords:
[…]
At the crossroads of applied AI, design, computer vision and robotics, this paper explores how a blimp could be used to non-disruptively analyze indoor crime scenes.
Crime is an important problem, ruining numerous lives: the global cost of violence has been estimated at a staggering $9.4 trillion USD, with homicide at $705.89 billion and assaults at $325.27 billion [1]. Furthermore, current approaches are not always successful, with low "clearance rates" the norm rather than the exception (e.g., 40% of homicides worldwide are not being solved, with worsening rates in various countries) [2]. As voiced in the UN’s 16th sustainable development goal (Peace and Justice), more needs to be done to prevent crime. However, crime often takes place behind closed doors and without witnesses, which requires careful documentation to unravel what happened. In such cases, "the crime scene investigator must set one goal above all others: secure the integrity of the crime scene . . . Lost or compromised evidence makes a crime scene investigator’s job harder and can seriously damage investigations. Blood spatter patterns and fingerprints can be inadvertently smudged, footprints or tire treads walked on if care is not taken, and trace evidence can be scattered hither and yon by those unaware of its very presence".111universalclass.com/articles/law/setting-crime-scene-perimeters.htm Evidence can also degrade naturally over time (e.g., DNA due to bacteria, heat, light, moisture, and mold) [3] when investigations are delayed due to under-staffing and under-funding–as in the saying "when seconds count, the police are only minutes away". To preserve indoor crime scenes more carefully and automatically, we imagine that blimps with cameras could be used to float over and record potentially sensitive evidence.
The term "blimp" refers to a non-rigid, lighter-than-air aircraft that possesses actuators and sensors within a "gondola" (or "nacelle") to enable flight and monitoring.222https://www.airships.net/dirigible Here we use the term in a relaxed sense, permitting some small rigid structure, in the belief that our ideas could be also generally relevant for other kinds of airships, zeppelins, lighter-than-air unmanned aerial vehicles (LTA-UAVs), dirigibles, balloons, and aerostats. While aerodynes like drones could also be useful for indoor crime scene analysis, we believe that blimps feature some merits such as safety, silence, stability (ability to hover over long periods of time), sustainability (low cost, low energy/emission), soft landing, easy transport when deflated, and/or minimal training/licensing requirements for operators [4]. Conversely, the demerits of blimps, such as slow speed, large size, and sensitivity to wind, might be less problematic indoors in smaller domestic areas that are quick to cover, while still wide enough for humans to move around, and potentially less windy than large outdoor sites.
A challenge is that such blimps intended to operate within difficult human environments in the context of crime would require a complex array of hardware and software capabilities. To tackle this challenge, we adopted a rapid prototyping methodology that emphasizes "flexibility, possibility, and design insights as incubated subjectively through the designer", seeking to "cast our net widely" [5]. Thus, our contribution lies in reporting on some opportunities and limitations revealed during our design experience.
The remainder of the paper is organized as follows: In Section 2, we summarize some previous work on blimps for security and indoor use. Details from our exploration in Section 3 are summarized in Section 4, which also points out some next waypoints on our journey to supporting safe and just societies.
2 Related Work
The basic promise of flying platforms for facilitating crime scene analysis has been explored, by manually piloting (or simulating flight of) a drone to spot mock-up bloodstains, guns, knives, and bodies [6, 7, 8]. We have also previously presented an overview of various tasks such a flying robot could perform at a crime scene [9]. However, Bucknell and Bassindale raised the alarm that downwash (wind from a drone’s propellers) could disturb sensitive evidence [10]. This could be problematic, since fibers and fine particles have been used, in some important but difficult cases, to catch various criminals, such as the "Route 40 Killer" or "Green River Killer".333https://listverse.com/2022/03/05/10-criminals-who-were-caught-thanks-to-trace-fiber-evidence/,444https://www.nbcnews.com/news/us-news/gary-ridgway-green-river-serial-killer-washington-rcna67794 In the end, Bucknell and Bassindale concluded that drones should be flown at high heights [10]. This was a consequence of how this investigation was conducted, which involved piloting a Parrot AR.Drone 2.0 at various heights. However, some wind could still be generated, and a drone might not be able to distance itself sufficiently when flying under low ceilings or detecting high-up evidence on a table or shelf. Thus, it seemed likely to us that less wind might be generated by a blimp slowly drifting or hovering overhead without using propellers, which is our focus here.
More generally, the use of blimps for security applications has been considered previously: For example, blimps could be used by the police to monitor and deter crime [11].555http://www.reuters.com/article/2011/01/16/us-crime-blimp-utah-idUSTRE70F1DJ20110116 Saiki et al. proposed the use of a blimp to monitor disasters, describing control equations for a 12 m prototype with a stereo camera and LiDAR [12]. As well, the Ukrainian Armed Forces are also allegedly using tethered aerostats with infrared sensing from Aerobavovna and Kvertus that can help with surveillance and communication, act as signal relays, and launch drones when an enemy drone is detected [13]. The same article points out that China and the United States have also proposed aerostats that can launch drone swarms–an interesting idea also for crime scene analysis.
Various companies have also focused on designing small indoor blimps for entertainment: For example, a patent from 1994 (US5429542A) proposes a helium-filled remote-controlled saucer toy, and various toys have been sold, such as the Megatech Blimp in the early 2000s, a fish-shaped blimp called Air Swimmers in 2011 and the NanoBlimp in 2013.666https://en.wikipedia.org/wiki/Air_Swimmer,777https://ohgizmo.com/nanoblimp-is-allegedly-worlds-smallest-rc-blimp As well, the company Festo released many imaginative creations, such as a floating manta ray called in 2007, 888https://www.youtube.com/watch?v=c3-wIICjAhE a jellyfish called the Festo AirJelly in 2008, 999https://www.youtube.com/watch?v=divLsTtA5vk a penguin called the AirPenguin in 2009, 101010https://www.youtube.com/watch?v=jPGgl5VH5go a moving origami-style blimp called the SmartInversion in 2012, 111111https://www.youtube.com/watch?v=BDiTicX8HRU a swarm of dancing spherical blimps with LEDs called the eMotionSpheres in 2014, 121212https://www.youtube.com/watch?v=5iqP1oPZ3Qw and a spherical blimp that can pick up a bottle and hand it to a person, called FreeMotionHandling, in 2016. 131313https://www.youtube.com/watch?v=dv6zQr1_C9Q
Papers have also explored the design of novel kinds of indoor blimps for various applications. For example, in 2012, we built a first flying humanoid robot prototype, Angel, intended to safely interact with people; in addition to exploring how to plan safe motions and convey functional or emotional meanings through flight, we also experimented with two flight mechanisms, with wings or propellers [14]. 141414https://www.youtube.com/watch?v=4vbQ2tMOjgI/ In 2015, St-Onge et al. also made a uniquely-shaped cubic flying blimp, for indoor flight at art shows [15]. As well, in 2019, Yao et al. built a blimp to safely follow people indoors, detecting faces and two hand gestures via an RGB camera, and communicating via visual flight patterns and an LED display [16]. Furthermore, in 2019, Ferdous et al. built a blimp with low-current components, including an IR sensor and IMU, to enable usage over many hours [17]. In 2020, Seguin et al. used a convolutional neural network (CNN) to estimate the horizontal position and heading of a blimp from RGB video, along with a LiDAR to estimate height [18]. Recently, in 2024, Pham et al. described a bio-inspired blimp with fins and an onboard inertial measurement unit (IMU) [19]. Additionally, in 2024, Huang et al. designed an interesting 36" blimp with an OpenMV camera that can conduct remote measurement of important vital signs such as heart rate, respiration, or blood pressure via remote photoplethysmography; the blimp detects people with YOLO, approaches via PID control, and transmits data from stably detected foreheads [20]. Furthermore, in 2025, Xu et al. described a fish-shaped flapping-wing blimp named Cuddle-Fish that users enjoyed touching [21]. While offering insights, none of these studies considered the unique challenges of crime scene analysis by a blimp.
With regard to functions for movement, documentation, and sensing, much prior work with blimps seems to have focused on control algorithms to deal with wind for particular embodiments [22]–suggesting the usefulness of checking how easy it is to control the blimp manually. As well, videos taken by a blimp’s camera would require time to watch; given that various previous work has explored the use of 3D mapping of crime scenes [23], it seemed useful if the video could be used to generate a 3D model that can be immediately inspected. Additionally, various other modalities could be used for documenting evidence, such as heat, sound, smell, or distances through solids (via radar, or WiFi signals). For example, thermal traces could reveal recent activity, like jettisoned or hidden weapons [24], or drugs.
Looking further into the future, it could also be useful to check the feasibility of automatically detecting and classifying evidence, and how a blimp could plan itself how to move around to acquire data. For example, we have previously used YOLO, to detect suspicious objects [25]. Such tools could be used to gain insight into an important kind of evidence, blood. Bloodstains are fundamentally divided into three main kinds: passive drops falling due to gravity, active blood (e.g., flying from a victim’s body or from a weapon), or transfer stains (e.g., smears from a bloody hand or shoe walking through blood).151515https://www.forensicsciencesimplified.org/blood/principles.html/ While Bergman et al. recently described using a CNN to discriminate between passive drip vs. active spatter bloodstains [26], we are not aware of work that has tackled classifying between all three classes (possibly due to the sensitivity and paucity of publicly available crime data).
As well, in addition to making sense of what it "sees", an autonomous blimp would also be required to itself plan how to move to acquire data. The problem of how a robot should move to cover an area is known as coverage path planning (CPP). An NP-hard problem related to the Traveling Salesman Problem, CPP is relevant to various applications, from vacuum cleaning, to lawn mowing, snow removal, search-and-rescue, and aerial or underwater terrain exploration [27]. For example, typical simplified algorithms for a vacuum cleaner robot include snaking (also known as zigzag, snake, or s-path algorithm–typically combined with a random walk), as well as random motion, and spiraling [28]. Snaking was found to be comparatively efficient for achieving high coverage faster, with less unneeded passes–at the cost of more turning and potential weakness to wheel or sensor errors that could occur in cheaper robots. In more complex scenarios where optimized algorithms are desired, the state of the art involves deep reinforcement learning (RL; e.g., for vacuum cleaners [29]). Thus, various functions could be useful, which invited exploration.
3 Methods
To start exploring potential manual and automated capabilities, a basic platform was first required. With the aim of facilitating replication, cost-effective, standard-sized, off-the-shelf components and freely available software tools were used, as described below.
3.1 Set-up
3.1.1 Hardware
Various electronics components where used, such as those summarized in Table 1.
| Component | Weight | Current |
| Raspberry Pi Zero 2 W | 10 g | 460mA161616https://andrejacobs.org/b24/electronics/raspberry-pi-zero-2-w-temperature-and-power-consumption/,171717https://picockpit.com/raspberry-pi/everything-about-raspberry-pi-zero-2-w/ |
| ESP32-Cam | 10 g | 180- 310mA181818https://docs.sunfounder.com/projects/galaxy-rvr/en/latest/hardware/cpn_esp_32_cam.html,191919https://blog.arducam.com/esp32-machine-vision-learning-guide/ |
| TF-Luna | 5 g | 70mA (x 3 = 210mA)202020https://www.waveshare.com/wiki/TF-Luna_LiDAR_Range_Sensor |
| MLX90640 | 3.5 g | 20mA212121https://www.adafruit.com/product/4407 |
| Lipo Shim | 2.6 g | (15uA quiescent)222222https://www.pishop.us/product/pimoroni-zero-lipo-shim/ |
Computing. To obtain and wirelessly communicate sensor data, we set up a small Raspberry Pi Zero 2 W minicomputer (6.5cm × 3.0cm) with a 1 GHz quad-core 64-bit Arm Cortex-A53 CPU, 512 MB of SDRAM, and 2.4GHz WiFi.232323https://www.raspberrypi.com/products/raspberry-pi-zero-2-w As well, to facilitate user control of the blimp, an external laptop comprising an Intel i7 CPU @ 2.60GHz with 16 GB RAM was used.
Sensors. RGB Camera. A downward-facing ESP32-Cam was used to stream SVGA at 30fps from its 2 Megapixel OV2640 camera. 242424https://www.espressif.com/en/products/modules/esp32 (Various alternatives exist, such as the OpenMV camera, which although costing more, features more onboard processing capabilities.)
Lidars. To detect distances in front, to the side, and below the blimp (e.g., to enable wall-following or mapping), three TF-Lunas were used. TF-Luna is a small (3.5 cm x 2.1 cm x 1.3 cm, weight 5g) Time-of-Flight (TOF) single-point LiDAR based on a 850nm Vertical Cavity Surface Emitting Laser (VCSEL), which detects distances within a range of 0.2 m to 8 m with 1 cm resolution and 6 cm accuracy.252525https://www.waveshare.com/wiki/TF-Luna_LiDAR_Range_Sensor Again, various useful alternatives existed, like the VL53l1X sensor with 4 m range, but the TF-Luna’s longer range felt like it could be useful in larger rooms or corridors.
Thermal camera. Along the way, to explore possibilities for other modalities (e.g., detecting warm evidence that has recently been touched by a user), we attached a MLX90640 board. MLX90640 is a small (2.5 cm x 1.8 cm x 1.6 cm, 3.5 g) 768 pixel thermal camera with 110 x 70 field of view composed of a 24 x 32 array of IR thermal sensors that detect from -40 C to 300 C with an accuracy of 2C at a frame rate of 2 Hz.262626https://www.melexis.com/en/product/mlx90640/far-infrared-thermal-sensor-array
Power. To power the microcomputer and sensors, a "LiPo Shim" was used, which can provide a maximum of 1.5 A continuous current from a 1 cell lithium polymer battery (a low battery warning occurs at 3.4 V, followed by automatic shutdown at 3.0 V).272727https://shop.pimoroni.com/products/lipo-shim With our two small batteries (1000 mAh or 700 mAh), operation was approximately 20 minutes. As well, a separate 3 V CR2 battery was used to power the actuators.
Actuation and Body. To achieve motion without "reinventing the wheel", we modified an existing toy, comprising three DC motors attached to propellers and a radio receiver/transmitter. The propellers allow left and right rotation, as well as upward, downward, forward, and backward translation. To control the blimp via our own program, we connected the transmitter to our laptop via some common NPN bipolar transistors (2N2222) and an Arduino Uno.282828https://www.arduino.cc To lift the gondola, three balloons were used, each with approximately 80 cm diameter, 40 cm height, and a lift capacity of approximately 70-80 g when full (thus, 200-250 g in total). (Given the weight of helium and air (0.18 kg/m3 vs. 1.29 kg/m3), and a lift of 1.11 kg by a cubic meter of helium, this suggests there was somewhat less than a tenth of a cubic meter in each balloon.)292929https://specialtygases.messergroup.com/documents/20136/3998363/Balloon%2BHelium.pdf
During our design exploration, one observation that stood out was how tricky it can be to design small blimps with many sensors: Size. First, we found with a miniature toy unit and regular sized latex balloon that virtually no payload could be accommodated and small balloons lost helium quickly. Next, we tried using two balloons with a larger toy unit, but even though each individual sensor could be fairly light, the weight of busses, cables, tape and boards to connect and affix sensors quickly added up. Furthermore, much current was required for the microcomputer, sensors, and actuators, which in turn called for including an appropriately large battery. As well, to avoid voltage spikes from the motors resetting the microcomputer, separate power supplies were used. Thus, in the spirit of easing such restrictions and enabling freer exploration, our latest version became quite large, with three balloons (total height approximately 150 m with gondola and ballast pouch). Awkwardness. Some initial attempts to use single and double sided tape and velcro to connect the gondola and balloons didn’t provide good results, and aligning bulky components seeking to rise or fall was difficult. The gondola and balloons sometimes split, with heavy tape and velcro adhering just to one surface and ripping away material, creating unclean surfaces. Previously, working with the "Angel" blimp, we had attached hooks to a table and each balloon, but this time we wished avoid adding such weight. Instead, individual balloons were temporarily stowed under tables and desks to immobilize them. When combining balloons, we tried at one point using a low ceiling (2.3 m) in a corridor. However, this was a public space with people frequently needing to pass, and once, after returning from a quick errand, we glimpsed a passerby whacking our blimp playfully, causing the balloons to break apart. Thereafter, we propped the blimp against the bottom of a projector in our lab, which had a high ceiling (3 m+). We tied strings through each balloon’s tail section (where helium is inserted) and holes in the projector shelf, as well as to chairs below. To connect the balloons more securely, we modified some wall hooks to attach capped carbon fiber rods running up along both sides of the blimp. Even then, the top balloon sometimes flew off, so we attached set screws to both the bottom and top of the blimp to secure the rods firmly. (While we could have used a bigger balloon instead of using multiple smaller balloons, we believe this modularity is useful for exploratory design, since we can shrink or expand freely, also without wasting helium.) As well, to prevent the gondola from falling off, we made a plastic base that was firmly taped to the bottom balloon and further secured with threads going to the side hooks; then bolts and nuts were used to attach and remove the gondola. Another challenge was that the sensors had to be below the propellers, so rods were used to suspend a "second" small gondola from the first, built from light cardboard. Additionally, ballast was used to achieve neutral buoyancy in air, but we noticed that recalibration was required whenever we switched out different-sized batteries (1000 mAh or 700 mAh). As well, when testing the propellers of the gondola on a table, we were surprised when even powering motors for a brief instant was even to cause the entire gondola to spin into our laptop, causing a propeller to break; thereafter we took care to firmly tape down the gondola whenever testing it.
3.1.2 Software
Software was also required. Server and client programs were written to send LiDAR, RGB, and thermal data, as well as control signals, between the gondola and an external laptop. TigerVNC was used to remotely develop the software on the Raspberry Pi.303030https://tigervnc.org/ As well, we built a user interface with color-coded functions in Python and OpenCV,313131https://opencv.org/ shown in Fig. 3. While we created Arduino code to control the blimp in various ways, to avoid potential problems if communication broke down, we ended up using commands that initiate short bursts rather than continuous motion.
One observation when developing software was how little information we could find on how to set up a reasonably realistic mock-up crime scene for testing; e.g., what kind of evidence is typical and where it is usually found. In addition to the challenge that crime classification schemes differ regionally, in line with the diversity of criminal law systems, Google search hits for random crime scene generation seemed to bring up text summaries of crimes intended to help stimulate writers (e.g., "Murder Scene Generator"323232https://perchance.org/znevive9bn). We could also not find any datasets online of crime scene sketches. One reason for the lack of crime data could be due to a view that such data are sensitive and private, for victims and their families, and could also spark copy-cat crimes. While one or two mock-up sketches were available (e.g., Galanakis et al. made available a scan of one mock-up crime scene [30], and our partners at the Swedish Police also shared a scan of a mock-up crime scene at NFC Villa), as the saying goes, no two crime scenes are the same. Thus, to get started, we built two simplified programs to simulate how evidence might be distributed at a crime scene, as shown in Fig. 4. The first general program allows a user to select a floor plan and a kind of crime (homicide, assault, burglary, or arson), then randomly computes which kinds of evidence are found and where. For initial floor plans, we used (1) a LiDAR scan of the National Forensics Center (NFC) villa, used by our partners at the Swedish Police to train, and (2) a virtual depiction of Halmstad Intelligent Home (HINT), a 50 m2 mock-up sensorized living environment at our university. All probability parameters can be assigned or tweaked by the user. The second program, intended for initial testing, simply creates a single "heap" of evidence, taking a list of evidence and using it to fill a grid. Large evidence comprising more than one cell is filled in first, then smaller evidence.
3.2 Exploring Possibilities for Indoor CSA Blimps
With the developed platform, based on the literature, we speculated that it could be interesting to check how much wind the blimp creates, how easy it is to control manually, how recordings of evidence can be further processed into 3D maps, and if thermal traces could be used. Furthermore, looking toward the future, we wished to check automatic detection, classification of bloodstains, and how a blimp could plan itself how to move around to acquire data. For evidence to use in testing, we gathered a photo of a firearm, two knives (one large and metallic, and one small and plastic), a pair of shoes, and three sheets with mock-up blood (red carmine dye–E120) to simulate passive, active, and transfer stains.
3.2.1 Disturbance
To test the degree to which sensitive evidence might be disturbed by a blimp, we placed a Mastech MS6252A digital anemometer on the ground beside the mock-up evidence. As a result, the wind sensor reading was 0.0 m/s even when passing the blimp even quite close to the evidence ( 20 cm). For comparison, passing a typical small drone, the DJI Ryze Tello, over the evidence at 1.2 m resulted in a wind speed of 0.6 to 0.8 m/s, which notably caused most of our evidence to scatter (viz., the sheets of paper containing mock-up blood and the photo of a firearm).
3.2.2 Sensing while drifting
Next, we checked on the blimp’s ability to move and sense evidence. The blimp was controlled manually to obtain 10 data samples. In each trial, a random ordering of the evidence was calculated using our Python program, then the physical evidence was manually placed by a human. As a result, on average, the blimp captured 5.2 of 7 objects (74.3%) in its view. The reason for missing some evidence sometimes was because the blimp was susceptible to drafts and small imbalances that sometimes caused it to veer.
3.2.3 3D Mapping
For easy viewing, we wished to explore converting video from the blimp to a 3D model. To test the waters, a video from the blimp was passed to the free version of 3DF Zephyr to create a 3D textured mesh. The result is shown in the Fig. 5(a). Some rough areas existed, e.g., on the mannequin "corpse" to the left-hand side of the image), since the free version of this software only uses the first 50 photos extracted from the video. (Due to this limitation, we also sped up the video to ensure the whole scene would be included.) However, we believe that the time required to generate a 3D map (approximately 3 minutes) was promising.
3.2.4 Thermal Sensing
In each trial controlling the blimp, we experimented with also touching a knife and the shoes directly beforehand. As a result, we noted that some warm areas could be seen in the thermal data, as in Fig. 5(b), but in general, the warmed objects (especially the smaller knife) seemed hard to distinguish due to our sensor’s low resolution. As well, a checkerboard pattern appears due to the blimp’s motion and the manner in which the MLX90640 sensor updates (in two passes, such that half of the pixels update slightly later). Thus, we believe that a higher resolution sensor with different update process would be required some day in the future for identifying recently-touched, non-metallic objects. (The MLX90640 could however still serve some other use, such as detecting investigators to follow, or hidden persons).
3.2.5 Object detection
It could save time if the blimp could be automated to perform important tasks like detection. This could then be highlighted in a 3D map in reports to investigators, or used to focus the blimp’s attention to more thoroughly investigate certain areas. As an initial check, we tried inputting an image taken by the blimp to a pretrained generic YOLO detection model (YOLOv8l)333333https://docs.ultralytics.com/ and ChatGPT.343434hhttps://chatgpt.com The first result in Fig. 5 (c) shows that while one knife is detected correctly, a bloody handprint is mistaken as a horse, and other objects are not detected; these errors could be due to low resolution or slightly dark lighting (approximately 300 lux in our lab when the image was taken). ChatGPT, however, correctly detected all objects. This seemed promising to us, and suggested using either an adapted version of YOLO with sufficient data for guns and blood, or a foundation model.
3.2.6 Bloodstain Classification
To explore classifying the three main kinds of bloodstains, we tried processing a photo of our evidence heap. First we color-picked red, then assessed the area and eccentricity of an ellipse fitted to detected contours of reasonable size. We estimated that transfer patterns from a hand might have high area, passive drops should have low eccentricity, and active drops high eccentricity, as in Fig. 5(d). This resulted in an classification accuracy of 74.1% (20 of 27 detected blood contours being correctly classified). Adding in false positives from color-picking (e.g., detected in the red tags on the shoes), the overall accuracy was 69.0% (20 of 29), which seemed promising for initial prototyping.
3.2.7 Coverage path planning for crime scenes
An automated blimp could move without piloting. Given that a full solution to this complex problem wouldn’t be possible within the scope of the current paper, we brainstormed to identify some variables of interest and formulate the general problem. We believe the blimp’s path planning strategy should depend on:
How precise blimp’s movements are–if imprecise, the blimp should stay farther away to avoid crashing, and more overlap might be required.
How accurate the blimp’s sensors are–if inaccurate, the blimp should stay closer to the evidence to achieve proper documentation.
How urgent the task is–if the blimp is needed elsewhere, a rapid path could be calculated; else in an important case, more time can be taken.
If the blimp should also gather evidence–then it should probably be close.
How much detail is desired–if the blimp should see under obstacles like tables and beds.
In general, however, we believe that indoor blimp crime scene analysis is characterized by the following requirements:
Some speed is desired, given that evidence can become contaminated or degraded over time.
Some overlap is desired for 3D mapping; much (unneeded) overlap however would be undesirable since excess motion could increase the risk of disrupted evidence.
While human domestic environments can be highly complex (e.g., ceiling obstacles like lamps, fans, alarms, or chandeliers, and ground obstacles like chairs, clothes, etc. and L-shaped rooms or hallways and walk-in closets), rooms tend to often be rectangular.
The blimp should try to restrict turning and vertical changes to avoid using up its battery. For example, the blimp can fly over low objects (e.g., to see on tables, beds)–although, compared to an outdoors scene, height is more constrained.
People in the room might intentionally or unintentionally obstruct the robot.
The blimp could be acting as part of a swarm.
The environment might be known (e.g., from floor plans) or unknown.
Based on this, both simplified and more complex solutions seemed possible. For simplicity, we propose that a snaking algorithm combined with a random walk and/or wall-following could be a starting point for simple testing, that is intuitive and easy to implement and carry out. For maximum detail, the blimp could calculate a path where it flies just above surfaces; else for efficiency, the blimp could calculate a space below any ceiling obstacles and above any ground obstacles. As well, the blimp could scan from high-up first, then return to any areas that seem interesting (e.g., where evidence has been discovered). Alternatively, a deep RL algorithm could be used to document in a human-like way (e.g., based on data of human investigators): to achieve high efficiency (reduce flight time, avoiding redundant re-coverage), improve map quality (ensure good feature overlap and loop closure, while avoiding poor viewing angles, low lighting, occlusions, shadows or reflective surfaces), and be room-agnostic (function successfully in unknown or specially designed rooms or buildings, while avoiding hand-engineering edge cases), and robot-agnostic (different blimps could use the same process to learn how to move well). A cost might be more difficulty in the implementation (complex equations, much training data required, tuning), and loss of explainability.
4 Discussion
In summary, we explored a new design idea, of using a blimp to record sensitive evidence at an indoor crime scene. To our best knowledge, our new prototype is the first small indoor blimp to carry a thermal camera, or three LiDARs, or a minicomputer (rather than a microcontroller), etc., and various hardware challenges were discussed. Some software was also developed to simulate crime scenes. As a result of our exploration, we found some support that drones can disturb sensitive evidence and confirmed our premise that blimps can keep wind to a minimum. As well, manually piloting was possible but imperfect (sensing 74.3% of evidence). A 3D map was generated within a few minutes using free software. Some recommendations for thermal sensing were given. Toward automating crime scene analysis, all objects in the blimp’s camera view were detected correctly by ChatGPT. As well, the three main kinds of bloodstain were detected and classified with an accuracy of 74.1% via color-picking and consideration of contour area and eccentricity. Finally, we discussed path-planning considerations and strategies. As well, a video and code have been made available.353535https://youtu.be/5ep9UeChn68,363636https://github.com/martincooney/Crime_Blimp
4.1 Limitations and Future Work
The current study is limited by its exploratory design approach, working with mock-up evidence in a lab. The blimp was large, with component weights unoptimized; surface mount/low power components (like a Raspberry Pi Pico instead of Zero 2 W), or Dist-YOLO in place of LiDARs [31], to reduce weight and size. As well, we used a remote laptop, which, although common in the related studies noted, would not be secure at a real crime scene, given that criminals could find a way to hack (intercept or jam) communications. Our crime scene generating software is currently based on guesses and should be trained with real data to enable more realistic outputs. To help blimp operators, control algorithms should be developed. Other applications of heat sensing, and other miniature thermal cameras, should be tested. A more comprehensive study should be done on identifying and detecting the various kinds of evidence that could be important. For classifying bloodstains, only a single image was used. A reasonable number of (real) training and test data should be used in conjunction with a more advanced, machine-learning approach to achieve better results. As well, real composites of various kinds of bloodstains should be analyzed to infer the story of what happened, and further analysis can reveal, e.g., if a criminal is right or left-handed, the type of weapon used, what made a transfer stain (e.g., a hand or a shoe), which wounds were first, if death was immediate or not, and the number of blows, etc. Furthermore, path planning should be implemented and tested on our prototype.
As well, our focus was limited to sensitive crime evidence, whereas remote-monitoring blimps could be useful for other inspections of sensitive materials or even for helping to reduce the number of investigators ambushed and killed at crime scenes. By investigating these and other related directions, our aim is that opportunities could be found to contribute to a peaceful and happy society for all.
Acknowledgements.
We gratefully acknowledge support from the Swedish Innovation Agency (Vinnova) for the project "AI-Powered Crime Scene Analysis" as well as advice from the Swedish Police Authority. The initial idea for this paper stemmed from a conversation with Mikael Lilja (Superintendent, Police Region Öst).References
- Hoeffler [2017] A. Hoeffler, What are the costs of violence?, Politics, Philosophy & Economics 16 (2017) 422–445.
- Sturup et al. [2015] J. Sturup, D. Karlberg, M. Kristiansson, Unsolved homicides in sweden: A population-based study of 264 homicides, Forensic science international 257 (2015) 106–113.
- Benson et al. [2003] B. L. Benson, W. David, P. A. May, K. Harries, A. M. Rosenthal, A. Lorie, M. L. D. Fridell, G. Fisher-Stewart, J. Pedro, T. M. Saavedra, et al., Without a trace? advances in detecting trace evidence, National Institute of Justice 249 (2003).
- Rozhok [2023] A. Rozhok, Modeling of Innovative Lighter-than-Air UAV for Logistics, Surveillance, and Rescue Operations, Ph.D. thesis, Università degli studi di Genova, 2023.
- Zamfirescu-Pereira et al. [2021] J. Zamfirescu-Pereira, D. Sirkin, D. Goedicke, R. Lc, N. Friedman, I. Mandel, N. Martelaro, W. Ju, Fake it to make it: Exploratory prototyping in hri, in: Companion of the 2021 ACM/IEEE international conference on human-robot interaction, 2021, pp. 19–28.
- Urbanová et al. [2017] P. Urbanová, M. Jurda, T. Vojtíšek, J. Krajsa, Using drone-mounted cameras for on-site body documentation: 3d mapping and active survey, Forensic science international 281 (2017) 52–62.
- Georgiou et al. [2022] A. Georgiou, P. Masters, S. Johnson, L. Feetham, Uav-assisted real-time evidence detection in outdoor crime scene investigations, Journal of forensic sciences 67 (2022) 1221–1232.
- Araujo et al. [2019] P. Araujo, J. Fontinele, L. Oliveira, Multi-perspective object detection for remote criminal analysis using drones, IEEE Geoscience and Remote Sensing Letters 17 (2019) 1283–1286.
- Cooney et al. [2025] M. Cooney, S. Ponrajan, F. Alonso-Fernandez, Nano drone-based indoor crime scene analysis, arXiv preprint arXiv:2502.21019 (2025).
- Bucknell and Bassindale [2017] A. Bucknell, T. Bassindale, An investigation into the effect of surveillance drones on textile evidence at crime scenes, Science & justice 57 (2017) 373–375.
- Murphy et al. [2013] L. W. Murphy, C. Calabrese, J. Stanley, C. Crump, The future of drones in america: Law enforcement and privacy considerations, ACLU Statement for the Record for a Senate Judiciary Committee Hearing, American Civil Liberties Union 20 (2013).
- Saiki et al. [2011] H. Saiki, T. Fukao, T. Urakubo, T. Kohno, A path following control method under wind disturbances for outdoor blimp robots, in: 2011 IEEE/SICE International Symposium on System Integration (SII), IEEE, 2011, pp. 978–984.
- Trevithick [2025] J. Trevithick, Balloon-launched drone to intercept long range kamikaze drones emerges in ukraine, The War Zone (TMZ) (2025). URL: https://bmva-archive.org.uk/bmvc/2017/papers/paper188/index.html.
- Cooney et al. [2012] M. Cooney, F. Zanlungo, S. Nishio, H. Ishiguro, Designing a flying humanoid robot (fhr): effects of flight on interactive communication, in: 2012 IEEE RO-MAN: The 21st IEEE International Symposium on Robot and Human Interactive Communication, IEEE, 2012, pp. 364–371.
- St-Onge et al. [2015] D. St-Onge, C. Gosselin, N. Reeves, Dynamic modelling and control of a cubic flying blimp using external motion capture, Proceedings of the Institution of Mechanical Engineers, Part I: Journal of Systems and Control Engineering 229 (2015) 970–982.
- Yao et al. [2019] N.-s. Yao, Q.-y. Tao, W.-y. Liu, Z. Liu, Y. Tian, P.-y. Wang, T. Li, F. Zhang, Autonomous flying blimp interaction with human in an indoor space, Frontiers of Information Technology & Electronic Engineering 20 (2019) 45–59.
- Ferdous et al. [2019] S. U. Ferdous, A. Mohammadi, S. Lakshmanan, Developing a low-cost autonomous blimp with a reduced number of actuators, in: Unmanned Systems Technology XXI, volume 11021, SPIE, 2019, pp. 73–80.
- Seguin et al. [2020] L. Seguin, J. Zheng, A. Li, Q. Tao, F. Zhang, A deep learning approach to localization for navigation on a miniature autonomous blimp, in: 2020 IEEE 16th International Conference on Control & Automation (ICCA), IEEE, 2020, pp. 1130–1136.
- Pham et al. [2024] H. Q. Pham, S. Singh, M. Garratt, S. Ravi, Controlling a bio-inspired miniature blimp using a depth sensing neural-network camera, Bioinspiration & Biomimetics 19 (2024) 024001.
- Huang et al. [2024] H.-W. Huang, J. Chen, P. Rupp, C. Ehmke, P. R. Chai, R. Dhar, I. Ballinger, G. Traverso, Cost-effective blimp for autonomous and continuous vital signs monitoring, in: 2024 IEEE International Conference on Advanced Intelligent Mechatronics (AIM), IEEE, 2024, pp. 1553–1559.
- Xu et al. [2025] M. Xu, J. Shao, Y. Ju, X. Shen, Q. Gao, W. Chen, Q. Zhang, Y. S. Pai, G. Barbareschi, M. Hoppe, et al., Cuddle-fish: Exploring a soft floating robot with flapping wings for physical interactions, arXiv preprint arXiv:2504.01293 (2025).
- Bhat et al. [2024] S. S. Bhat, S. G. Anavatti, M. Garratt, S. Ravi, Review of autonomous outdoor blimps and their applications, Drone Systems and Applications 12 (2024) 1–21.
- Esposito et al. [2023] M. Esposito, F. Sessa, G. Cocimano, P. Zuccarello, S. Roccuzzo, M. Salerno, Advances in technologies in crime scene investigation. diagnostics 2023, 13, 3169. h ps, doi. org/10.3390/diagnostics13203169 Academic Editor: Hiroshi Ikegaya Received 15 (2023).
- Muñoz et al. [2025] J. D. Muñoz, J. Ruiz-Santaquiteria, O. Deniz, G. Bueno, Concealed weapon detection using thermal cameras, Journal of Imaging 11 (2025) 72.
- Cooney et al. [2024] M. Cooney, L. M. W. Klasén, F. Alonso-Fernandez, Designing robots to help women, in: 14th Scandinavian Conference on Artificial Intelligence (SCAI 2024): AI for a better society, June 10-11, 2024, Jönköping, Sweden, IEEE, 2024, pp. 168–177.
- Bergman et al. [2022] T. Bergman, M. Klöden, J. Dreßler, D. Labudde, Automatic classification of bloodstains with deep learning methods, KI-Künstliche Intelligenz 36 (2022) 135–141.
- Jonnarth et al. [2023] A. Jonnarth, J. Zhao, M. Felsberg, Learning coverage paths in unknown environments with deep reinforcement learning, arXiv preprint arXiv:2306.16978 (2023).
- Sörme and Edwards [2018] J. Sörme, T. Edwards, A comparison of path planning algorithms for robotic vacuum cleaners, 2018.
- Moon et al. [2022] W. Moon, B. Park, S. H. Nengroo, T. Kim, D. Har, Path planning of cleaning robot with reinforcement learning, in: 2022 IEEE International Symposium on Robotic and Sensors Environments (ROSE), IEEE, 2022, pp. 1–7.
- Galanakis et al. [2021] G. Galanakis, X. Zabulis, T. Evdaimon, S.-E. Fikenscher, S. Allertseder, T. Tsikrika, S. Vrochidis, A study of 3d digitisation modalities for crime scene investigation, Forensic sciences 1 (2021) 56–85.
- Vajgl et al. [2022] M. Vajgl, P. Hurtik, T. Nejezchleba, Dist-yolo: Fast object detection with distance estimation, Applied Sciences 12 (2022). URL: https://www.mdpi.com/2076-3417/12/3/1354. doi:10.3390/app12031354.

### v2

@@
 Abstract
-To tackle the crucial problem of crime, evidence at indoor crime scenes must be analyzed before it becomes contaminated or degraded.
-Here, as an application of artificial intelligence (AI), computer vision, and robotics, we explore how a blimp could be designed as a kind of "floating camera" to drift over and record evidence with minimal disturbance.
-In particular, rapid prototyping is used to develop a proof-of-concept to gain insight into what such blimps could do, manually piloted or semi-autonomously.
-As a result, we show the feasibility of attaching various components to an indoor blimp, and confirm our basic premise, that blimps can sense evidence without producing much wind.
-Some additional suggestions–regarding mapping, sensing, and path-finding–aim to stimulate the flow of ideas for further exploration.
+Crime is a critical problem—which often takes place behind closed doors, posing additional difficulties for investigators.
+To bring hidden truths to light, evidence at indoor crime scenes must be documented before any contamination or degradation occurs.
+Here, we address this challenge from the perspective of artificial intelligence (AI), computer vision, and robotics: Specifically, we explore the use of a blimp as a "floating camera" to drift over and record evidence with minimal disturbance.
+Adopting a rapid prototyping approach, we develop a proof-of-concept to investigate capabilities required for manual or semi-autonomous operation.
+Consequently, our results demonstrate the feasibility of equipping indoor blimps with various components (such as RGB and thermal cameras, LiDARs, and WiFi, with 20 minutes of battery life).
+Moreover, we confirm the core premise: that such blimps can be used to observe crime scene evidence while generating little airflow.
+We conclude by proposing some ideas related to detection (e.g., of bloodstains), mapping, and path planning, with the aim of stimulating further discussion and exploration.
 keywords:
@@
 At the crossroads of applied AI, design, computer vision and robotics, this paper explores how a blimp could be used to non-disruptively analyze indoor crime scenes.
-Crime is an important problem, ruining numerous lives: the global cost of violence has been estimated at a staggering $9.4 trillion USD, with homicide at $705.89 billion and assaults at $325.27 billion [1].
-Furthermore, current approaches are not always successful, with low "clearance rates" the norm rather than the exception (e.g., 40% of homicides worldwide are not being solved, with worsening rates in various countries) [2].
-As voiced in the UN’s 16th sustainable development goal (Peace and Justice), more needs to be done to prevent crime.
-However, crime often takes place behind closed doors and without witnesses, which requires careful documentation to unravel what happened.
-In such cases, "the crime scene investigator must set one goal above all others: secure the integrity of the crime scene .
+Crime remains a global crisis, ruining numerous lives and draining public resources: Intentional homicides cruelly took approximately half of a million lives in 2021—the highest rate in twenty years [1].
+The global cost of violence has also been estimated at a staggering $9.4 trillion USD, with homicide at $705.89 billion and assaults at $325.27 billion [2].
+Furthermore, current countermeasures are often unsuccessful, with 40% of homicides worldwide going unsolved, and worsening clearance rates noted in countries such as the United States (US), Canada, Trinidad, and Tobago [3].
+As emphasized in the United Nation’s 16th sustainable development goal (Peace and Justice), new approaches are urgently needed to improve crime prevention and investigation [1].
+Yet, crime often takes place indoors—behind closed doors and without witnesses—requiring careful, often painstaking documentation to piece together what transpired.
+In such cases, as noted in forensic literature, "the crime scene investigator must set one goal above all others: secure the integrity of the crime scene .
 .
@@
 Lost or compromised evidence makes a crime scene investigator’s job harder and can seriously damage investigations.
-Blood spatter patterns and fingerprints can be inadvertently smudged, footprints or tire treads walked on if care is not taken, and trace evidence can be scattered hither and yon by those unaware of its very presence".111universalclass.com/articles/law/setting-crime-scene-perimeters.htm Evidence can also degrade naturally over time (e.g., DNA due to bacteria, heat, light, moisture, and mold) [3] when investigations are delayed due to under-staffing and under-funding–as in the saying "when seconds count, the police are only minutes away".
-To preserve indoor crime scenes more carefully and automatically, we imagine that blimps with cameras could be used to float over and record potentially sensitive evidence.
-The term "blimp" refers to a non-rigid, lighter-than-air aircraft that possesses actuators and sensors within a "gondola" (or "nacelle") to enable flight and monitoring.222https://www.airships.net/dirigible Here we use the term in a relaxed sense, permitting some small rigid structure, in the belief that our ideas could be also generally relevant for other kinds of airships, zeppelins, lighter-than-air unmanned aerial vehicles (LTA-UAVs), dirigibles, balloons, and aerostats.
-While aerodynes like drones could also be useful for indoor crime scene analysis, we believe that blimps feature some merits such as safety, silence, stability (ability to hover over long periods of time), sustainability (low cost, low energy/emission), soft landing, easy transport when deflated, and/or minimal training/licensing requirements for operators [4].
-Conversely, the demerits of blimps, such as slow speed, large size, and sensitivity to wind, might be less problematic indoors in smaller domestic areas that are quick to cover, while still wide enough for humans to move around, and potentially less windy than large outdoor sites.
-A challenge is that such blimps intended to operate within difficult human environments in the context of crime would require a complex array of hardware and software capabilities.
+Blood spatter patterns and fingerprints can be inadvertently smudged, footprints or tire treads walked on if care is not taken, and trace evidence can be scattered hither and yon by those unaware of its very presence".111https://universalclass.com/articles/law/setting-crime-scene-perimeters.htm Furthermore, time is a critical factor: Evidence can degrade due to bacteria, heat, light, moisture, or mold [4], and investigations are often delayed or cut short by understaffing or limited resources.
+As the saying goes, "when seconds count, the police are only minutes away." Thus, to help preserve indoor crime scenes more carefully and automatically, we imagine that blimps with cameras could be used to float over and record potentially sensitive evidence.
+The term "blimp" traditionally refers to a non-rigid, lighter-than-air aircraft equipped with actuators and sensors within a “gondola” or “nacelle” to enable flight and monitoring.222https://www.airships.net/dirigible Here, we adopt the term in a relaxed sense—e.g., allowing for some small rigid structure—in the belief that our ideas could be also broadly relevant for other kinds of aerostats, airships, dirigibles, and lighter-than-air unmanned aerial vehicles (LTA-UAVs), like hybrid dynastats or rotastats, zeppelins, and balloons.
+While aerodynes such as drones could also aid in indoor crime scene analysis, blimps offer some unique advantages: safety, silence, stability (ability to hover over extended periods), sustainability (due to low cost, energy use, and emissions), soft landing, easy transport when deflated, and minimal training or licensing requirements for operators [5].
+Conversely, typical disadvantages—such as sensitivity to wind, low speed, and large size—could be less limiting in indoor environments, which are generally less windy and more compact, yet wide enough to be navigable by humans.
+A challenge is that blimps intended to help solve difficult crimes in chaotic human environments would require a complex array of hardware and software capabilities.
 To tackle this challenge, we adopted a rapid prototyping methodology that emphasizes "flexibility, possibility, and design insights as incubated subjectively through the designer", seeking to "cast our net widely" [6].
@@
 2 Related Work
-The basic promise of flying platforms for facilitating crime scene analysis has been explored, by manually piloting (or simulating flight of) a drone to spot mock-up bloodstains, guns, knives, and bodies [6, 7, 8].
-We have also previously presented an overview of various tasks such a flying robot could perform at a crime scene [9].
-However, Bucknell and Bassindale raised the alarm that downwash (wind from a drone’s propellers) could disturb sensitive evidence [10].
-This could be problematic, since fibers and fine particles have been used, in some important but difficult cases, to catch various criminals, such as the "Route 40 Killer" or "Green River Killer".333https://listverse.com/2022/03/05/10-criminals-who-were-caught-thanks-to-trace-fiber-evidence/,444https://www.nbcnews.com/news/us-news/gary-ridgway-green-river-serial-killer-washington-rcna67794 In the end, Bucknell and Bassindale concluded that drones should be flown at high heights [10].
-This was a consequence of how this investigation was conducted, which involved piloting a Parrot AR.Drone 2.0 at various heights.
-However, some wind could still be generated, and a drone might not be able to distance itself sufficiently when flying under low ceilings or detecting high-up evidence on a table or shelf.
-Thus, it seemed likely to us that less wind might be generated by a blimp slowly drifting or hovering overhead without using propellers, which is our focus here.
-More generally, the use of blimps for security applications has been considered previously: For example, blimps could be used by the police to monitor and deter crime [11].555http://www.reuters.com/article/2011/01/16/us-crime-blimp-utah-idUSTRE70F1DJ20110116 Saiki et al.
-proposed the use of a blimp to monitor disasters, describing control equations for a 12 m prototype with a stereo camera and LiDAR [12].
-As well, the Ukrainian Armed Forces are also allegedly using tethered aerostats with infrared sensing from Aerobavovna and Kvertus that can help with surveillance and communication, act as signal relays, and launch drones when an enemy drone is detected [13].
-The same article points out that China and the United States have also proposed aerostats that can launch drone swarms–an interesting idea also for crime scene analysis.
-Various companies have also focused on designing small indoor blimps for entertainment: For example, a patent from 1994 (US5429542A) proposes a helium-filled remote-controlled saucer toy, and various toys have been sold, such as the Megatech Blimp in the early 2000s, a fish-shaped blimp called Air Swimmers in 2011 and the NanoBlimp in 2013.666https://en.wikipedia.org/wiki/Air_Swimmer,777https://ohgizmo.com/nanoblimp-is-allegedly-worlds-smallest-rc-blimp As well, the company Festo released many imaginative creations, such as a floating manta ray called in 2007, 888https://www.youtube.com/watch?v=c3-wIICjAhE a jellyfish called the Festo AirJelly in 2008, 999https://www.youtube.com/watch?v=divLsTtA5vk a penguin called the AirPenguin in 2009, 101010https://www.youtube.com/watch?v=jPGgl5VH5go a moving origami-style blimp called the SmartInversion in 2012, 111111https://www.youtube.com/watch?v=BDiTicX8HRU a swarm of dancing spherical blimps with LEDs called the eMotionSpheres in 2014, 121212https://www.youtube.com/watch?v=5iqP1oPZ3Qw and a spherical blimp that can pick up a bottle and hand it to a person, called FreeMotionHandling, in 2016.
-131313https://www.youtube.com/watch?v=dv6zQr1_C9Q
-Papers have also explored the design of novel kinds of indoor blimps for various applications.
-For example, in 2012, we built a first flying humanoid robot prototype, Angel, intended to safely interact with people; in addition to exploring how to plan safe motions and convey functional or emotional meanings through flight, we also experimented with two flight mechanisms, with wings or propellers [14].
-141414https://www.youtube.com/watch?v=4vbQ2tMOjgI/ In 2015, St-Onge et al.
-also made a uniquely-shaped cubic flying blimp, for indoor flight at art shows [15].
+The basic potential of flying platforms for facilitating crime scene analysis has been explored, by manually piloting drones—or simulating their flight—to spot mock-up bloodstains, guns, knives, and bodies [7, 8, 9].
+We have also previously presented an overview of tasks that a flying robot could perform at a crime scene [10].
+However, Bucknell and Bassindale raised the alarm that downwash—wind from a drone’s propellers—could disturb sensitive evidence [11].
+This could be problematic, since fibers and fine particles have played a crucial role in solving some important but challenging cases, such as those of the "Route 40" or "Green River" killers.333https://listverse.com/2022/03/05/10-criminals-who-were-caught-thanks-to-trace-fiber-evidence/,444https://www.nbcnews.com/news/us-news/gary-ridgway-green-river-serial-killer-washington-rcna67794 In their study, Bucknell and Bassindale concluded that drones should be flown at higher altitudes, based on experiments involving a Parrot AR.Drone 2.0 at varying heights [11].
+Nonetheless, wind could still be generated, especially in indoor settings with low ceilings or when examining elevated evidence on tables or shelves.
+In contrast, a blimp that hovers or drifts slowly without using its propellers could generate less airflow—an assumption that motivates the present study.
+More broadly, blimps have been previously considered for security-related applications: For instance, Murphy and colleagues described how the police could use blimps to monitor and deter crime—an idea discussed in the city of Ogden in the US [12].555https://www.blimpinfo.com/uncategorized/plans-for-ogden-police-blimp-canceled/ Saiki et al.
+also proposed the use of a blimp to monitor disasters, describing control equations for a 12 m prototype with a stereo camera and LiDAR [13].
+As well, the Ukrainian Armed Forces are using tethered aerostats with infrared sensing from Aerobavovna and Kvertus to help with surveillance, communication, and drone deployment when an enemy drone is detected [14].
+The same article notes that China and the US have also proposed aerostats that can launch drone swarms—an idea that could also hold promise for analysis of larger crime scenes.
+Various companies have also designed small indoor blimps, for entertainment: For example, a 1994 patent (US5429542A) proposed a helium-filled, remote-controlled saucer toy; over the years, various toys have followed, such as the Megatech Blimp in the early 2000s, a fish-shaped blimp called Air Swimmers in 2011 and the NanoBlimp in 2013.666https://en.wikipedia.org/wiki/Air_Swimmer,777https://ohgizmo.com/nanoblimp-is-allegedly-worlds-smallest-rc-blimp As well, the company Festo has stirred interest with many imaginative creations: e.g., Air_ray, a floating manta ray introduced in 2007;888https://www.youtube.com/watch?v=c3-wIICjAhE AirJelly, a jellyfish from 2008;999https://www.youtube.com/watch?v=divLsTtA5vk AirPenguin, a penguin from 2009;101010https://www.youtube.com/watch?v=jPGgl5VH5go SmartInversion, a moving origami-style blimp from 2012;111111https://www.youtube.com/watch?v=BDiTicX8HRU eMotionSpheres, a swarm of dancing spherical blimps with LEDs from 2014;121212https://www.youtube.com/watch?v=5iqP1oPZ3Qw and FreeMotionHandling, a spherical blimp that can pick up a bottle and hand it to a person, from 2016.131313https://www.youtube.com/watch?v=dv6zQr1_C9Q
+Several research papers have also explored the design of novel indoor blimps for diverse applications: For example, in 2012, we built a first flying humanoid robot prototype, Angel, intended to safely interact with people [15].141414https://www.youtube.com/watch?v=4vbQ2tMOjgI/ In addition to exploring how to plan safe motions and convey functional or emotional meanings through flight, we also experimented with two flight mechanisms, utilizing wings or propellers.
+In 2015, St-Onge et al.
+made a uniquely-shaped cubic flying blimp, for indoor flight at art shows [16].
 As well, in 2019, Yao et al.
 built a blimp to safely follow people indoors, detecting faces and two hand gestures via an RGB camera, and communicating via visual flight patterns and an LED display [17].
-Furthermore, in 2019, Ferdous et al.
-built a blimp with low-current components, including an IR sensor and IMU, to enable usage over many hours [17].
+That same year, Ferdous et al.
+built a blimp with low-current components, including an infrared (IR) sensor and inertial measurement unit (IMU), to enable usage over many hours [18].
 In 2020, Seguin et al.
@@
 Recently, in 2024, Pham et al.
-described a bio-inspired blimp with fins and an onboard inertial measurement unit (IMU) [19].
+described a bio-inspired blimp with fins and an onboard IMU [20].
 Additionally, in 2024, Huang et al.
@@
 Furthermore, in 2025, Xu et al.
-described a fish-shaped flapping-wing blimp named Cuddle-Fish that users enjoyed touching [21].
-While offering insights, none of these studies considered the unique challenges of crime scene analysis by a blimp.
-With regard to functions for movement, documentation, and sensing, much prior work with blimps seems to have focused on control algorithms to deal with wind for particular embodiments [22]–suggesting the usefulness of checking how easy it is to control the blimp manually.
-As well, videos taken by a blimp’s camera would require time to watch; given that various previous work has explored the use of 3D mapping of crime scenes [23], it seemed useful if the video could be used to generate a 3D model that can be immediately inspected.
-Additionally, various other modalities could be used for documenting evidence, such as heat, sound, smell, or distances through solids (via radar, or WiFi signals).
-For example, thermal traces could reveal recent activity, like jettisoned or hidden weapons [24], or drugs.
-Looking further into the future, it could also be useful to check the feasibility of automatically detecting and classifying evidence, and how a blimp could plan itself how to move around to acquire data.
+described a fish-shaped, flapping-wing blimp named Cuddle-Fish that users enjoyed touching [22].
+The same year, Hong and Tanaka presented unique walking humanoids with balloon torsos and articulated legs containing LEDs and various sensors, that can be controlled with gamepads or directed via external fans at an art installation [23].
+While informative, none of these studies considered the unique challenges of crime scene analysis by a blimp.
+With regard to functions for movement, documentation, and sensing to enable manual piloting, much prior work on blimps seems to have focused on control algorithms to deal with wind for particular embodiments [24]—suggesting the usefulness of checking how easy it is to pilot a CSA blimp manually.
+As well, videos captured by a blimp’s camera require time to watch; given that various previous work has explored the 3D mapping of crime scenes [25], it seemed useful if a video could be leveraged to generate a 3D model that can be immediately inspected.
+Additionally, various other modalities could be used for documenting evidence, such as heat, sound, smell, or distances through solids (via radar or WiFi signals).
+For example, thermal traces could reveal recent activity, like jettisoned or hidden weapons [26] or drugs.
+Looking further ahead, it could also be beneficial to explore the feasibility of automatically detecting and classifying evidence, and how a blimp itself could plan to move to acquire data.
 For example, we have previously used YOLO, to detect suspicious objects [27].
-Such tools could be used to gain insight into an important kind of evidence, blood.
-Bloodstains are fundamentally divided into three main kinds: passive drops falling due to gravity, active blood (e.g., flying from a victim’s body or from a weapon), or transfer stains (e.g., smears from a bloody hand or shoe walking through blood).151515https://www.forensicsciencesimplified.org/blood/principles.html/ While Bergman et al.
+Such a tool could also be used to gain insight into an important kind of evidence, blood.
+Bloodstains are generally categorized into three main kinds: passive (e.g., from drops falling under gravity), active (e.g., from blood expelled from a victim’s body or a weapon), or transfer (e.g., from contact by a bloody hand or shoe).151515https://www.forensicsciencesimplified.org/blood/principles.html While Bergman et al.
 recently described using a CNN to discriminate between passive drip vs.
-active spatter bloodstains [26], we are not aware of work that has tackled classifying between all three classes (possibly due to the sensitivity and paucity of publicly available crime data).
-As well, in addition to making sense of what it "sees", an autonomous blimp would also be required to itself plan how to move to acquire data.
-The problem of how a robot should move to cover an area is known as coverage path planning (CPP).
-An NP-hard problem related to the Traveling Salesman Problem, CPP is relevant to various applications, from vacuum cleaning, to lawn mowing, snow removal, search-and-rescue, and aerial or underwater terrain exploration [27].
-For example, typical simplified algorithms for a vacuum cleaner robot include snaking (also known as zigzag, snake, or s-path algorithm–typically combined with a random walk), as well as random motion, and spiraling [28].
-Snaking was found to be comparatively efficient for achieving high coverage faster, with less unneeded passes–at the cost of more turning and potential weakness to wheel or sensor errors that could occur in cheaper robots.
-In more complex scenarios where optimized algorithms are desired, the state of the art involves deep reinforcement learning (RL; e.g., for vacuum cleaners [29]).
-Thus, various functions could be useful, which invited exploration.
+active spatter bloodstains [28], we are not aware of work that has tackled classification across all three categories—possibly due to the sensitivity and scarcity of publicly available crime data.
+In addition to interpreting what it "sees", an autonomous blimp should also plan how to move to acquire data.
+The problem of how a robot should move to cover a given area is known as coverage path planning (CPP).
+An NP-hard problem related to the Traveling Salesman Problem, CPP arises in a wide range of applications, from vacuum cleaning, to lawn mowing, snow removal, search-and-rescue, and aerial or underwater terrain exploration [29].
+For example, typical simplified algorithms used by robotic vacuum cleaners include snaking (also known as a zigzag or s-path algorithm—often combined with random walks), as well as spiraling and random motion [30].
+Among these, snaking was found to be efficient at achieving high coverage faster, with fewer unneeded passes—at the cost of increased turning and weakness to wheel or sensor errors that can occur in lower-cost robots.
+In more complex scenarios where optimized algorithms are desired, the state of the art increasingly involves deep reinforcement learning (RL; e.g., for vacuum cleaners [31]).
+Accordingly, various manual and autonomous capabilities appeared like they could be useful, inviting exploration.
 3 Methods
-To start exploring potential manual and automated capabilities, a basic platform was first required.
+To start exploring potential capabilities, a basic platform was first required.
 With the aim of facilitating replication, cost-effective, standard-sized, off-the-shelf components and freely available software tools were used, as described below.
@@
 | MLX90640 | 3.5 g | 20 mA212121https://www.adafruit.com/product/4407 |
-| Lipo Shim | 2.6 g | (15uA quiescent)222222https://www.pishop.us/product/pimoroni-zero-lipo-shim/ |
+| Lipo Shim | 2.6 g | (15 A quiescent)222222https://www.pishop.us/product/pimoroni-zero-lipo-shim/ |
 Computing.
-To obtain and wirelessly communicate sensor data, we set up a small Raspberry Pi Zero 2 W minicomputer (6.5cm × 3.0cm) with a 1 GHz quad-core 64-bit Arm Cortex-A53 CPU, 512 MB of SDRAM, and 2.4GHz WiFi.232323https://www.raspberrypi.com/products/raspberry-pi-zero-2-w As well, to facilitate user control of the blimp, an external laptop comprising an Intel i7 CPU @ 2.60GHz with 16 GB RAM was used.
+To obtain and wirelessly communicate sensor data, we set up a small Raspberry Pi Zero 2 W minicomputer (6.5 cm × 3.0 cm) with a 1 GHz quad-core 64-bit Arm Cortex-A53 CPU, 512 MB of SDRAM, and 2.4 GHz WiFi.232323https://www.raspberrypi.com/products/raspberry-pi-zero-2-w As well, to control the blimp, an external laptop comprising an Intel i7 CPU @ 2.60GHz with 16 GB RAM and a 6GB GPU was used.
 Sensors.
 RGB Camera.
-A downward-facing ESP32-Cam was used to stream SVGA at 30fps from its 2 Megapixel OV2640 camera.
-242424https://www.espressif.com/en/products/modules/esp32 (Various alternatives exist, such as the OpenMV camera, which although costing more, features more onboard processing capabilities.)
+A downward-facing ESP32-Cam was used to stream SVGA at 30fps from its 2 Megapixel OV2640 camera.242424https://www.espressif.com/en/products/modules/esp32 (Various alternatives exist, such as the OpenMV camera,252525https://openmv.io which although costing more, features more onboard processing capabilities.)
 LiDARs.
-To detect distances in front, to the side, and below the blimp (e.g., to enable wall-following or mapping), three TF-Lunas were used.
-TF-Luna is a small (3.5 cm x 2.1 cm x 1.3 cm, weight 5g) Time-of-Flight (TOF) single-point LiDAR based on a 850nm Vertical Cavity Surface Emitting Laser (VCSEL), which detects distances within a range of 0.2 m to 8 m with 1 cm resolution and 6 cm accuracy.252525https://www.waveshare.com/wiki/TF-Luna_LiDAR_Range_Sensor Again, various useful alternatives existed, like the VL53l1X sensor with 4 m range, but the TF-Luna’s longer range felt like it could be useful in larger rooms or corridors.
+To detect distances in front, to the side, and below the blimp (e.g., allowing wall-following or mapping), three TF-Lunas were used.
+TF-Luna is a small (3.5 cm x 2.1 cm x 1.3 cm, weight 5g) Time-of-Flight (TOF) single-point LiDAR based on an 850 nm Vertical Cavity Surface Emitting Laser (VCSEL), which detects distances within a range of 0.2 m to 8 m with 1 cm resolution and 6 cm accuracy.262626https://www.waveshare.com/wiki/TF-Luna_LiDAR_Range_Sensor Again, alternatives existed, like the VL53l1X sensor with 4 m range—yet the TF-Luna’s longer range felt like it could be useful in larger rooms or corridors.
 Thermal camera.
-Along the way, to explore possibilities for other modalities (e.g., detecting warm evidence that has recently been touched by a user), we attached a MLX90640 board.
-MLX90640 is a small (2.5 cm x 1.8 cm x 1.6 cm, 3.5 g) 768 pixel thermal camera with 110 x 70 field of view composed of a 24 x 32 array of IR thermal sensors that detect from -40 C to 300 C with an accuracy of 2C at a frame rate of 2 Hz.262626https://www.melexis.com/en/product/mlx90640/far-infrared-thermal-sensor-array
+Along the way, to explore possibilities for other modalities (e.g., detecting warm evidence that has recently been touched by a user), we attached an MLX90640 board.
+MLX90640 is a small (2.5 cm x 1.8 cm x 1.6 cm, 3.5 g) 768 pixel thermal camera with a 110 x 70 field of view, composed of a 24 x 32 array of IR thermal sensors that detect from -40C to 300C with an accuracy of 2C at a frame rate of 2 Hz.272727https://www.melexis.com/en/product/mlx90640/far-infrared-thermal-sensor-array More powerful but costly alternatives such as FLIR exist.
 Power.
-To power the microcomputer and sensors, a "LiPo Shim" was used, which can provide a maximum of 1.5 A continuous current from a 1 cell lithium polymer battery (a low battery warning occurs at 3.4 V, followed by automatic shutdown at 3.0 V).272727https://shop.pimoroni.com/products/lipo-shim With our two small batteries (1000 mAh or 700 mAh), operation was approximately 20 minutes.
+To power the microcomputer and sensors, a "LiPo Shim" was used, which provides a maximum of 1.5 A continuous current from a 1 cell lithium polymer battery (a low battery warning occurs at 3.4 V, followed by automatic shutdown at 3.0 V).282828https://shop.pimoroni.com/products/lipo-shim With our two small batteries (700 mAh or 1000 mAh, approx.
+15-20 g), operation was approximately 20 minutes.
 As well, a separate 3 V CR2 battery was used to power the actuators.
@@
 The propellers allow left and right rotation, as well as upward, downward, forward, and backward translation.
-To control the blimp via our own program, we connected the transmitter to our laptop via some common NPN bipolar transistors (2N2222) and an Arduino Uno.282828https://www.arduino.cc To lift the gondola, three balloons were used, each with approximately 80 cm diameter, 40 cm height, and a lift capacity of approximately 70-80 g when full (thus, 200-250 g in total).
+To control the blimp via our own program, we connected the transmitter to our laptop via some common NPN bipolar transistors (2N2222) and an Arduino Uno.292929https://www.arduino.cc To lift the gondola, three balloons were used, each with approximately 80 cm diameter, 40 cm height, and a lift capacity of approximately 70-80 g when full (thus, 200-250 g in total).
 (Given the weight of helium and air (0.18 kg/m3 vs.
-1.29 kg/m3), and a lift of 1.11 kg by a cubic meter of helium, this suggests there was somewhat less than a tenth of a cubic meter in each balloon.)292929https://specialtygases.messergroup.com/documents/20136/3998363/Balloon%2BHelium.pdf
-During our design exploration, one observation that stood out was how tricky it can be to design small blimps with many sensors: Size.
-First, we found with a miniature toy unit and regular sized latex balloon that virtually no payload could be accommodated and small balloons lost helium quickly.
-Next, we tried using two balloons with a larger toy unit, but even though each individual sensor could be fairly light, the weight of busses, cables, tape and boards to connect and affix sensors quickly added up.
-Furthermore, much current was required for the microcomputer, sensors, and actuators, which in turn called for including an appropriately large battery.
-As well, to avoid voltage spikes from the motors resetting the microcomputer, separate power supplies were used.
-Thus, in the spirit of easing such restrictions and enabling freer exploration, our latest version became quite large, with three balloons (total height approximately 150 m with gondola and ballast pouch).
-Awkwardness.
-Some initial attempts to use single and double sided tape and velcro to connect the gondola and balloons didn’t provide good results, and aligning bulky components seeking to rise or fall was difficult.
-The gondola and balloons sometimes split, with heavy tape and velcro adhering just to one surface and ripping away material, creating unclean surfaces.
-Previously, working with the "Angel" blimp, we had attached hooks to a table and each balloon, but this time we wished avoid adding such weight.
-Instead, individual balloons were temporarily stowed under tables and desks to immobilize them.
-When combining balloons, we tried at one point using a low ceiling (2.3 m) in a corridor.
-However, this was a public space with people frequently needing to pass, and once, after returning from a quick errand, we glimpsed a passerby whacking our blimp playfully, causing the balloons to break apart.
-Thereafter, we propped the blimp against the bottom of a projector in our lab, which had a high ceiling (3 m+).
-We tied strings through each balloon’s tail section (where helium is inserted) and holes in the projector shelf, as well as to chairs below.
-To connect the balloons more securely, we modified some wall hooks to attach capped carbon fiber rods running up along both sides of the blimp.
-Even then, the top balloon sometimes flew off, so we attached set screws to both the bottom and top of the blimp to secure the rods firmly.
-(While we could have used a bigger balloon instead of using multiple smaller balloons, we believe this modularity is useful for exploratory design, since we can shrink or expand freely, also without wasting helium.) As well, to prevent the gondola from falling off, we made a plastic base that was firmly taped to the bottom balloon and further secured with threads going to the side hooks; then bolts and nuts were used to attach and remove the gondola.
-Another challenge was that the sensors had to be below the propellers, so rods were used to suspend a "second" small gondola from the first, built from light cardboard.
-Additionally, ballast was used to achieve neutral buoyancy in air, but we noticed that recalibration was required whenever we switched out different-sized batteries (1000 mAh or 700 mAh).
-As well, when testing the propellers of the gondola on a table, we were surprised when even powering motors for a brief instant was even to cause the entire gondola to spin into our laptop, causing a propeller to break; thereafter we took care to firmly tape down the gondola whenever testing it.
+1.29 kg/m3), and a lift of 1.11 kg by a cubic meter of helium, this suggests there was somewhat less than a tenth of a cubic meter in each balloon.)303030https://specialtygases.messergroup.com/balloon-helium
+During our design exploration, we observed that equipping small blimps with many sensors posed challenges: Size.
+Miniature toy units with regular latex balloons offered minimal payload capacity and lost helium quickly.
+Using two balloons with a larger toy improved lift, but the cumulative weight of sensors, boards, cables, tape, and power systems added up fast.
+High current demands from the microcomputer, sensors, and actuators required a large battery, and separate power supplies were adopted to avoid voltage spikes that could reset the microcontroller.
+To overcome these constraints and enable free exploration, our latest version became substantially larger, using three balloons (total height 150 cm with gondola and ballast).
+Assembly.
+Initial attachment methods using tape or velcro proved unreliable; gondolas detached or tore balloon surfaces.
+Unlike our earlier “Angel” design, where balloons were anchored to a table, we avoided heavy hooks this time by temporarily tucking balloons under tables to immobilize them.
+At one point, we used a 2.3 m corridor ceiling to assemble the blimp, but public access made it impractical—once, we glimpsed a passerby whacking our blimp playfully, causing the balloons to break apart.
+We later secured the blimp under a ceiling-mounted projector in our lab, tying balloon tails also to chairs as needed, and connecting balloons via capped carbon fiber rods held by modified wall hooks.
+Despite this, the top balloon sometimes detached, so we added set screws at both ends to secure the rods.
+To prevent the gondola from detaching, we taped a plastic base to the bottom balloon, reinforced with threads to side hooks, and affixed the gondola using bolts and nuts.
+Sensors had to be mounted below the propellers, so we suspended a second lightweight cardboard gondola from the main one using rods.
+We also used ballast to fine-tune buoyancy, but noticed that battery changes (e.g., between 700 mAh and 1000 mAh) required recalibration.
+While testing propellers on a tabletop, we were also surprised when only briefly powering a motor caused the entire gondola to spin and hit a laptop, breaking a propeller—after which we made sure to tape down the gondola securely when testing.
 3.1.2 Software
-Software was also required.
-Server and client programs were written to send LiDAR, RGB, and thermal data, as well as control signals, between the gondola and an external laptop.
-TigerVNC was used to remotely develop the software on the Raspberry Pi.303030https://tigervnc.org/ As well, we built a user interface with color-coded functions in Python and OpenCV,313131https://opencv.org/ shown in Fig.
+We also developed server and client programs to transmit LiDAR, RGB, and thermal data, as well as control signals between the gondola and our external laptop.
+TigerVNC was used to remotely develop the software on the Raspberry Pi.313131https://tigervnc.org/ As well, we built a user interface with color-coded controls in Python and OpenCV,323232https://opencv.org/ shown in Fig.
 3.
-While we created Arduino code to control the blimp in various ways, to avoid potential problems if communication broke down, we ended up using commands that initiate short bursts rather than continuous motion.
-One observation when developing software was how little information we could find on how to set up a reasonably realistic mock-up crime scene for testing; e.g., what kind of evidence is typical and where it is usually found.
-In addition to the challenge that crime classification schemes differ regionally, in line with the diversity of criminal law systems, Google search hits for random crime scene generation seemed to bring up text summaries of crimes intended to help stimulate writers (e.g., "Murder Scene Generator"323232https://perchance.org/znevive9bn).
-We could also not find any datasets online of crime scene sketches.
-One reason for the lack of crime data could be due to a view that such data are sensitive and private, for victims and their families, and could also spark copy-cat crimes.
-While one or two mock-up sketches were available (e.g., Galanakis et al.
-made available a scan of one mock-up crime scene [30], and our partners at the Swedish Police also shared a scan of a mock-up crime scene at NFC Villa), as the saying goes, no two crime scenes are the same.
-Thus, to get started, we built two simplified programs to simulate how evidence might be distributed at a crime scene, as shown in Fig.
-4.
-The first general program allows a user to select a floor plan and a kind of crime (homicide, assault, burglary, or arson), then randomly computes which kinds of evidence are found and where.
-For initial floor plans, we used (1) a LiDAR scan of the National Forensics Center (NFC) villa, used by our partners at the Swedish Police to train, and (2) a virtual depiction of Halmstad Intelligent Home (HINT), a 50 m2 mock-up sensorized living environment at our university.
-All probability parameters can be assigned or tweaked by the user.
-The second program, intended for initial testing, simply creates a single "heap" of evidence, taking a list of evidence and using it to fill a grid.
-Large evidence comprising more than one cell is filled in first, then smaller evidence.
+Various Arduino functions were implemented to allow the blimp to move, but to reduce risks if communication was lost, we ultimately chose to use commands that initiate short bursts rather than continuous motion.
+A key challenge to our exploration was a lack of information on how to set up reasonably realistic mock-up crime scenes: e.g., what kind of evidence is typical and where it is usually found.
+We found no publicly available datasets of crime scene layouts—likely due to privacy, legal, or safety concerns, or regional variances on how to categorize crimes.
+Aside from two scans of mock-up crime scene shared by Galanakis et al.
+[32] and the Swedish Police from their National Forensics Center (NFC) training facility, and a few mock-up sketches on the internet, most search results seemed to point to story-writing tools (e.g., “Murder Scene Generator”).333333https://perchance.org/znevive9bn This was problematic, since, as the saying goes, "no two crime scenes are the same", blimps should be able to operate in different environments, and engineering evaluations typically require more than two or three trials (deep learning systems can even require thousands of training samples).
+Thus, to support initial testing, we created two simplified simulation tools.
+The first lets users choose a crime type (e.g., homicide, assault, burglary, or arson) and floor plan (e.g., NFC Villa or our university’s 50 m2 "HINT" smart home), then randomly generates evidence types and locations, using customizable probabilities.
+The second tool places all evidence in a single heap, filling a grid by size—larger items first.
 3.2 Exploring Possibilities for Indoor CSA Blimps
-With the developed platform, based on the literature, we speculated that it could be interesting to check how much wind the blimp creates, how easy it is to control manually, how recordings of evidence can be further processed into 3D maps, and if thermal traces could be used.
-Furthermore, looking toward the future, we wished to check automatic detection, classification of bloodstains, and how a blimp could plan itself how to move around to acquire data.
-For evidence to use in testing, we gathered a photo of a firearm, two knives (one large and metallic, and one small and plastic), a pair of shoes, and three sheets with mock-up blood (red carmine dye–E120) to simulate passive, active, and transfer stains.
+The developed platform was applied to check the questions identified in reviewing the literature: how much wind the blimp creates, how easy it is to control manually, how recordings of evidence can be further processed into 3D maps, and if thermal traces could be used.
+Looking ahead, we also aimed to explore the feasibility of automatic detection, classification of bloodstains, and autonomous path planning by the blimp for data acquisition.
+For evidence to use in testing, we gathered a photo of a firearm, two knives (one large and metallic, the other small and plastic), a pair of shoes, and three sheets with mock-up blood (red carmine dye—E120) to simulate passive, active, and transfer stains.
 3.2.1 Disturbance
 To test the degree to which sensitive evidence might be disturbed by a blimp, we placed a Mastech MS6252A digital anemometer on the ground beside the mock-up evidence.
-As a result, the wind sensor reading was 0.0 m/s even when passing the blimp even quite close to the evidence ( 20 cm).
-For comparison, passing a typical small drone, the DJI Ryze Tello, over the evidence at 1.2 m resulted in a wind speed of 0.6 to 0.8 m/s, which notably caused most of our evidence to scatter (viz., the sheets of paper containing mock-up blood and the photo of a firearm).
+As a result, the wind sensor consistently read 0.0 m/s, even when the blimp passed close (20 cm) to the evidence.
+For comparison, flying a typical small drone—the DJI Ryze Tello—over the evidence at 1.2 m resulted in wind speeds of 0.6 to 0.8 m/s, which notably caused most of our evidence to scatter (viz., the sheets of paper containing mock-up blood and the photo of a firearm).
 3.2.2 Sensing while Drifting
 Next, we checked on the blimp’s ability to move and sense evidence.
-The blimp was controlled manually to obtain 10 data samples.
+The blimp was manually controlled to obtain data ten times.
 In each trial, a random ordering of the evidence was calculated using our Python program, then the physical evidence was manually placed by a human.
 As a result, on average, the blimp captured 5.2 of 7 objects (74.3%) in its view.
-The reason for missing some evidence sometimes was because the blimp was susceptible to drafts and small imbalances that sometimes caused it to veer.
+Some objects were occasionally missed because the blimp was susceptible to drafts and small imbalances that sometimes caused it to veer.
 3.2.3 3D Mapping
-For easy viewing, we wished to explore converting video from the blimp to a 3D model.
-To test the waters, a video from the blimp was passed to the free version of 3DF Zephyr to create a 3D textured mesh.
-The result is shown in the Fig.
+For easy viewing, we explored converting video from the blimp to a 3D model.
+To test the waters, a video from the blimp was processed using the free version of 3DF Zephyr to create a 3D textured mesh.343434https://www.3dflow.net/3df-zephyr-photogrammetry-software/ The result is shown in Fig.
 5(a).
-Some rough areas existed, e.g., on the mannequin "corpse" to the left-hand side of the image), since the free version of this software only uses the first 50 photos extracted from the video.
-(Due to this limitation, we also sped up the video to ensure the whole scene would be included.) However, we believe that the time required to generate a 3D map (approximately 3 minutes) was promising.
+Some areas appeared rough—for example, around the mannequin "corpse" on the left side of the image—since the free software version only processes the first 50 frames extracted from the video.
+(Due to this limitation, we also sped up the video to ensure the entire scene was included.) However, the short time required to generate a 3D map (approximately 3 minutes) seemed promising.
 3.2.4 Thermal Sensing
-In each trial controlling the blimp, we experimented with also touching a knife and the shoes directly beforehand.
-As a result, we noted that some warm areas could be seen in the thermal data, as in Fig.
-5(b), but in general, the warmed objects (especially the smaller knife) seemed hard to distinguish due to our sensor’s low resolution.
-As well, a checkerboard pattern appears due to the blimp’s motion and the manner in which the MLX90640 sensor updates (in two passes, such that half of the pixels update slightly later).
-Thus, we believe that a higher resolution sensor with different update process would be required some day in the future for identifying recently-touched, non-metallic objects.
-(The MLX90640 could however still serve some other use, such as detecting investigators to follow, or hidden persons).
+We also experimented with briefly touching a knife and the shoes before each trial controlling the blimp.
+As a result, some warm areas were visible in the thermal data, as exemplified in Fig.
+5(b).
+However, in general, the heated objects—especially the smaller knife—seemed hard to distinguish due to the low resolution of our thermal sensor.
+As well, a checkerboard pattern appeared due to the blimp’s motion and the manner in which the MLX90640 sensor updates (in two passes, such that half of the pixels update slightly later).
+Thus, we believe that a higher-resolution sensor with a different update mechanism would be required to identify recently touched, non-metallic objects.
+That said, the MLX90640 could still be useful for other purposes, such as detecting investigators (to follow) or hidden persons.
 3.2.5 Object Detection
-It could save time if the blimp could be automated to perform important tasks like detection.
-This could then be highlighted in a 3D map in reports to investigators, or used to focus the blimp’s attention to more thoroughly investigate certain areas.
-As an initial check, we tried inputting an image taken by the blimp to a pretrained generic YOLO detection model (YOLOv8l)333333https://docs.ultralytics.com/ and ChatGPT.343434hhttps://chatgpt.com The first result in Fig.
-5 (c) shows that while one knife is detected correctly, a bloody handprint is mistaken as a horse, and other objects are not detected; these errors could be due to low resolution or slightly dark lighting (approximately 300 lux in our lab when the image was taken).
-ChatGPT, however, correctly detected all objects.
-This seemed promising to us, and suggested using either an adapted version of YOLO with sufficient data for guns and blood, or a foundation model.
+Time and effort could be saved by automating key tasks such as detection of evidence.
+Detection results could also be highlighted in a 3D map in reports to investigators, or used to focus the blimp’s attention on areas requiring closer inspection.
+As an initial check, we tried passing an image taken by the blimp to a pretrained generic YOLOv8l detection model353535https://docs.ultralytics.com/ and ChatGPT.363636https://chatgpt.com As shown in Fig.
+5(c), the YOLO model correctly identified one knife, but a bloody handprint was mistaken as a horse, and other objects were not detected.
+These errors were likely due to low image resolution or dim lighting conditions (approximately 300 lux in our lab when the image was taken).
+ChatGPT, by contrast, correctly detected all objects.
+This seemed promising and suggested the usefulness of applying a generic foundation model, if hardware permits and a specific model fine-tuned on a sufficient number of images of firearms and blood is unavailable.
 3.2.6 Bloodstain Classification
-To explore classifying the three main kinds of bloodstains, we tried processing a photo of our evidence heap.
-First we color-picked red, then assessed the area and eccentricity of an ellipse fitted to detected contours of reasonable size.
-We estimated that transfer patterns from a hand might have high area, passive drops should have low eccentricity, and active drops high eccentricity, as in Fig.
+To explore classifying the three main kinds of bloodstains, we tried processing a photo of our evidence "heap".
+First, we used color-picking and contours to identify clusters of red pixels.
+Then, we assessed the area and eccentricity of an ellipse fitted to each detected contour of reasonable size.
+We assumed that typical transfer stains (e.g., from a hand or foot) could be large compared to individual droplets, that passive drops should appear round (exhibit low eccentricity), and that active drops should appear elongated (have high eccentricity), as illustrated in Fig.
 5(d).
-This resulted in an classification accuracy of 74.1% (20 of 27 detected blood contours being correctly classified).
-Adding in false positives from color-picking (e.g., detected in the red tags on the shoes), the overall accuracy was 69.0% (20 of 29), which seemed promising for initial prototyping.
+This resulted in a classification accuracy of 74.1%, with 20 of 27 detected blood contours correctly classified.
+Adding in false positives from color-picking (e.g., red tags on the shoes), the overall accuracy dropped to 69.0% (20 of 29), which still seemed promising for initial prototyping.
 3.2.7 Coverage Path Planning for Crime Scenes
-An automated blimp could move without piloting.
-Given that a full solution to this complex problem wouldn’t be possible within the scope of the current paper, we brainstormed to identify some variables of interest and formulate the general problem.
-We believe the blimp’s path planning strategy should depend on:
-How precise blimp’s movements are–if imprecise, the blimp should stay farther away to avoid crashing, and more overlap might be required.
-How accurate the blimp’s sensors are–if inaccurate, the blimp should stay closer to the evidence to achieve proper documentation.
-How urgent the task is–if the blimp is needed elsewhere, a rapid path could be calculated; else in an important case, more time can be taken.
-If the blimp should also gather evidence–then it should probably be close.
-How much detail is desired–if the blimp should see under obstacles like tables and beds.
-In general, however, we believe that indoor blimp crime scene analysis is characterized by the following requirements:
-Some speed is desired, given that evidence can become contaminated or degraded over time.
-Some overlap is desired for 3D mapping; much (unneeded) overlap however would be undesirable since excess motion could increase the risk of disrupted evidence.
-While human domestic environments can be highly complex (e.g., ceiling obstacles like lamps, fans, alarms, or chandeliers, and ground obstacles like chairs, clothes, etc.
-and L-shaped rooms or hallways and walk-in closets), rooms tend to often be rectangular.
-The blimp should try to restrict turning and vertical changes to avoid using up its battery.
-For example, the blimp can fly over low objects (e.g., to see on tables, beds)–although, compared to an outdoors scene, height is more constrained.
-People in the room might intentionally or unintentionally obstruct the robot.
-The blimp could be acting as part of a swarm.
-The environment might be known (e.g., from floor plans) or unknown.
-Based on this, both simplified and more complex solutions seemed possible.
-For simplicity, we propose that a snaking algorithm combined with a random walk and/or wall-following could be a starting point for simple testing, that is intuitive and easy to implement and carry out.
-For maximum detail, the blimp could calculate a path where it flies just above surfaces; else for efficiency, the blimp could calculate a space below any ceiling obstacles and above any ground obstacles.
-As well, the blimp could scan from high-up first, then return to any areas that seem interesting (e.g., where evidence has been discovered).
-Alternatively, a deep RL algorithm could be used to document in a human-like way (e.g., based on data of human investigators): to achieve high efficiency (reduce flight time, avoiding redundant re-coverage), improve map quality (ensure good feature overlap and loop closure, while avoiding poor viewing angles, low lighting, occlusions, shadows or reflective surfaces), and be room-agnostic (function successfully in unknown or specially designed rooms or buildings, while avoiding hand-engineering edge cases), and robot-agnostic (different blimps could use the same process to learn how to move well).
-A cost might be more difficulty in the implementation (complex equations, much training data required, tuning), and loss of explainability.
+An automated blimp could operate without manual piloting.
+Given that a full solution to the complex problem of autonomous coverage path planning for a CSA blimp would not be possible within the scope of the current paper, we constrained our brainstorming here to identifying key variables, requirements and potential strategies.
+First, we propose that designs should consider the following factors:
+Blimp’s capabilities
+Movement precision: If the blimp’s movements are imprecise, it should fly at a safer, farther distance to avoid collisions and plan more overlap to ensure coverage.
+Sensor accuracy: If sensors are less accurate, the blimp should fly closer to the evidence to capture sufficient detail.
+Task
+Task urgency: If time is constrained and the blimp is needed elsewhere, speed and efficiency can be prioritized.
+Else, in critical cases, more thorough scanning could be appropriate.
+For example, higher-detail documentation could require maneuvering beneath or around obstacles like tables and beds.
+Task specification: If the blimp is expected to also gather evidence, it should move closer to objects of interest.
+Thus, in general, we believe that indoor crime scene mapping by a blimp should meet the following requirements, listed in arbitrary order:
+Speed: Timely mapping is desired, prior to contamination or degradation of evidence, also given battery life constraints (e.g., mapping a room should take an hour or less).
+Overlap: Some overlap (e.g., 25%) is required for 3D reconstruction, but excessive motion could disturb the scene.
+Navigability: While some environments could be mostly empty and of simple rectangular shape, advanced capabilities could be required to navigate in complex domestic spaces with obstacles such as ceiling fans, lamps, chairs, or hallways, even if floor plans are known.
+Efficiency: The blimp should minimize turning and vertical movement to conserve battery life.
+Flying at a greater height than the tallest standing object could be efficient, although indoor spaces constrain vertical movement more than outdoor ones.
+Interaction: Humans, animals, or dynamic obstacles such as other robots could unintentionally or intentionally obstruct the blimp. Safety is paramount.
+Adaptability: An advanced blimp could operate as part of a swarm, use an LLM to prioritize regions likely to contain evidence, and remain within a geofenced region to avoid disturbing people or objects.
+Given these requirements, both simple and advanced planning approaches seemed possible:
+Basic approaches: A snaking strategy combined with random walks, wall-following, or behavior-based robotics, at a collision-free height, could serve as a practical and intuitive baseline for early-stage testing.
+Variable-height approaches: For maximum detail, the blimp could calculate a close path just above each surface.
+Conversely, initial high-altitude scans could be followed by targeted low-altitude re-visits to interesting areas (e.g., where potential evidence has been detected).
+AI-driven approaches: Deep reinforcement learning (RL) could be used to mimic how investigators explore crime scenes.
+This could achieve high efficiency (minimizing flight time and avoiding redundant re-coverage), map quality (ensuring good feature overlap and loop closure, while avoiding poor viewing angles, low lighting, occlusions, shadows or reflective surfaces), room-agnosticism (functioning successfully in unknown or specially designed rooms or buildings, while avoiding hand-engineering edge cases), and robot-agnosticism (allowing different blimps to use the same process to learn how to move well).
+However, complex methods—such as deep RL—could also introduce challenges, including difficult implementation, large data requirements, extensive tuning requirements, and reduced explainability.
 4 Discussion
-In summary, we explored a new design idea, of using a blimp to record sensitive evidence at an indoor crime scene.
-To our best knowledge, our new prototype is the first small indoor blimp to carry a thermal camera, or three LiDARs, or a minicomputer (rather than a microcontroller), etc., and various hardware challenges were discussed.
-Some software was also developed to simulate crime scenes.
-As a result of our exploration, we found some support that drones can disturb sensitive evidence and confirmed our premise that blimps can keep wind to a minimum.
-As well, manually piloting was possible but imperfect (sensing 74.3% of evidence).
-A 3D map was generated within a few minutes using free software.
-Some recommendations for thermal sensing were given.
-Toward automating crime scene analysis, all objects in the blimp’s camera view were detected correctly by ChatGPT.
+In this paper, we explored a new design idea, of using a blimp to document sensitive evidence at an indoor crime scene.
+Proof of Concept (Hardware and Software).
+To our knowledge, our new prototype is the first small indoor blimp to carry a thermal camera, three LiDARs, or a minicomputer (rather than a microcontroller).
+In addition to discussing various hardware challenges, some software was also developed to pilot the blimp and "generate" crime scenes.
+Manual operation.
+As a result of our exploration, we found some further support that drones can disturb sensitive evidence and confirmed our premise that blimps can keep wind to a minimum.
+As well, manually piloting was possible but imperfect, with an average of 74.3% of objects captured in view in a single pass.
+Furthermore, a 3D map was generated within a few minutes using free software, and some recommendations for thermal sensing were given.
+Towards Automation.
+All objects in the blimp’s camera view were detected correctly by ChatGPT.
 As well, the three main kinds of bloodstain were detected and classified with an accuracy of 74.1% via color-picking and consideration of contour area and eccentricity.
 Finally, we discussed path-planning considerations and strategies.
-As well, a video and code have been made available.353535https://youtu.be/5ep9UeChn68,363636https://github.com/martincooney/Crime_Blimp
+As well, a video and code have been made available.373737https://youtu.be/5ep9UeChn68,383838https://github.com/martincooney/Crime_Blimp
 4.1 Limitations and Future Work
-The current study is limited by its exploratory design approach, working with mock-up evidence in a lab.
-The blimp was large, with component weights unoptimized; surface mount/low power components (like a Raspberry Pi Pico instead of Zero 2 W), or Dist-YOLO in place of LiDARs [31], to reduce weight and size.
-As well, we used a remote laptop, which, although common in the related studies noted, would not be secure at a real crime scene, given that criminals could find a way to hack (intercept or jam) communications.
-Our crime scene generating software is currently based on guesses and should be trained with real data to enable more realistic outputs.
-To help blimp operators, control algorithms should be developed.
-Other applications of heat sensing, and other miniature thermal cameras, should be tested.
-A more comprehensive study should be done on identifying and detecting the various kinds of evidence that could be important.
-For classifying bloodstains, only a single image was used.
-A reasonable number of (real) training and test data should be used in conjunction with a more advanced, machine-learning approach to achieve better results.
-As well, real composites of various kinds of bloodstains should be analyzed to infer the story of what happened, and further analysis can reveal, e.g., if a criminal is right or left-handed, the type of weapon used, what made a transfer stain (e.g., a hand or a shoe), which wounds were first, if death was immediate or not, and the number of blows, etc.
-Furthermore, path planning should be implemented and tested on our prototype.
-As well, our focus was limited to sensitive crime evidence, whereas remote-monitoring blimps could be useful for other inspections of sensitive materials or even for helping to reduce the number of investigators ambushed and killed at crime scenes.
-By investigating these and other related directions, our aim is that opportunities could be found to contribute to a peaceful and happy society for all.
+The current study is limited by its exploratory design approach, working with mock-up evidence in a lab:
+Unoptimized components and guesswork.
+To reduce size and weight, surface mount or low power components (e.g., Raspberry Pi Pico instead of Zero 2 W), Dist-YOLO in place of LiDARs [33], or smaller batteries could be evaluated.
+Hybrid setups could also allow greater speed and payloads.
+As well, our current reliance on a remote laptop—though common in related work—would pose security risks at a real crime scene, given that criminals could find a way to hack (intercept or jam) communications.
+Thus, dedicated onboard hardware acceleration for computer vision is desired.
+For basic tasks, an NPU (neural processing unit) could suffice, such as the 1 GHz unit on OpenMV N6 that can run YOLO at 30 FPS.393939https://openmv.io/collections/all-products/products/openmv-n6 Heavier Foundation Models or LLMs requiring GPUs (Graphic Processor Units) with sufficient VRAM (Video Random Access Memory) could also become feasible on future blimp platforms; e.g., 16 GB for a 30B LLM with 4-16 bit quantization.404040https://pyvideotrans.com/en/_posts/deployllm Our crime scene generation software is also heuristic-based and would benefit from training with real data to enable more realistic outputs.
+Simplified approaches.
+Control algorithms should be developed to help blimp operators.
+Given the low thrust of blimp propellers, we should clarify if drifting might not even be required in some contexts.
+Additional use cases for thermal sensing and alternative miniature thermal cameras should be explored, as well as other free 3D Mapping tools (that allow more than fifty photos).
+Operation in challenging conditions—such as darkness or smoke—should also be explored, for example using auxiliary lighting, low-light cameras, or near-infrared imaging.
+Ultimately, field evaluations with real crime scenes and data will be essential to develop real products and services.
+Limited data and testing.
+A more comprehensive study is required to identify the various kinds of evidence that could be important and how they can be detected.
+Also, for classifying bloodstains, only a single image was used; a reasonably-sized dataset—ideally real—should be used to train and evaluate a more advanced machine learning model, to achieve better results.
+Furthermore, in-depth forensic interpretation of bloodstain composites could reveal valuable information such as weapon type, handedness, number of blows, sequence of injuries, sources of transfer stains (e.g., a hand or a shoe), or immediacy of death.
+Path planning should also be implemented and tested; for this, an omnidirectional camera could also improve scene coverage while minimizing unnecessary flight, reducing the risk of disturbing sensitive evidence.
+Beyond documentation of evidence at indoor crime scenes, remote-controlled blimps could even aid in hostage or standoff situations by offering silent surveillance to inform response planning, and reduce investigator exposure to ambushes or danger.
+Blimps could also benefit other sensitive applications where airflow or noise from conventional drones could be detrimental—for instance, munitions or chemical facilities with ignition risks, nuclear disaster sites where radioactive dust should not be stirred, avalanche or building sites in danger of collapsing, habitats of endangered animals sensitive to noise, or pharmaceutical cleanrooms or semiconductor plants where slight air movements could introduce flaws.
+By pursuing such directions, our aim is to identify opportunities for technology to contribute to a more peaceful and happy society for all.
 Acknowledgements.
-We gratefully acknowledge support from the Swedish Innovation Agency (Vinnova) for the project "AI-Powered Crime Scene Analysis" as well as advice from the Swedish Police Authority.
-The initial idea for this paper stemmed from a conversation with Mikael Lilja (Superintendent, Police Region Öst).References
+We gratefully acknowledge support from the Swedish Innovation Agency (Vinnova) for the project "AI-Powered Crime Scene Analysis" as well as various advice from the Swedish Police Authority (the initial idea for this paper stemmed from a conversation with Mikael Lilja).References
+- un2 [2023] The Sustainable Development Goals Report 2023, 2023. URL: https://unstats.un.org/sdgs/report/2023/The-Sustainable-Development-Goals-Report-2023.pdf, accessed: 2025-05-26.
 - Hoeffler [2017] A. Hoeffler, What are the costs of violence?, Politics, Philosophy & Economics 16 (2017) 422–445.
@@
 Oliveira, Multi-perspective object detection for remote criminal analysis using drones, IEEE Geoscience and Remote Sensing Letters 17 (2019) 1283–1286.
-- Cooney et al. [2025] M. Cooney, S. Ponrajan, F. Alonso-Fernandez, Nano drone-based indoor crime scene analysis, arXiv preprint arXiv:2502.21019 (2025).
+- Cooney et al.
+[2025] M.
+Cooney, S.
+Ponrajan, F.
+Alonso-Fernandez, Nano Drone-based Indoor Crime Scene Analysis, 2025 IEEE International Conference on Advanced Robotics and its Social Impacts (ARSO) (accepted April 29, 2025; arXiv preprint arXiv:2502.21019) (2025).
 - Bucknell and Bassindale [2017] A. Bucknell, T. Bassindale, An investigation into the effect of surveillance drones on textile evidence at crime scenes, Science & justice 57 (2017) 373–375.
@@
 Hoppe, et al., Cuddle-Fish: Exploring a Soft Floating Robot with Flapping Wings for Physical Interactions, arXiv preprint arXiv:2504.01293 (2025).
+- Hong and Tanaka [2025] D.
+Hong, Y.
+Tanaka, Buoyant choreographies: Harmonies of light, sound, and human connection, in: IEEE International Conference on Robotics and Automation (ICRA) 2025, IEEE, 2025.
 - Bhat et al. [2024] S. S. Bhat, S. G. Anavatti, M. Garratt, S. Ravi, Review of autonomous outdoor blimps and their applications, Drone Systems and Applications 12 (2024) 1–21.
@@
 Roccuzzo, M.
-Salerno, Advances in technologies in crime scene investigation.
-diagnostics 2023, 13, 3169.
-h ps, doi.
-org/10.3390/diagnostics13203169 Academic Editor: Hiroshi Ikegaya Received 15 (2023).
+Salerno, Advances in Technologies in Crime Scene Investigation, Diagnostics 2023, 13(20), 3169.
+DOI org/10.3390/diagnostics13203169 (2023).
 - Muñoz et al. [2025] J. D. Muñoz, J. Ruiz-Santaquiteria, O. Deniz, G. Bueno, Concealed Weapon Detection Using Thermal Cameras, Journal of Imaging 11 (2025) 72.